OpenStudy (anonymous):

A rectangle has an area of x^2-y^2 square units. What is the width of the rectangle

OpenStudy (solomonzelman):

Well, they are probably looking for something like \(\normalsize\color{blue}{ x^2-y^2=(x+y)~~(x-y) }\) \(\normalsize\color{blue}{ ↓ }\) \(\normalsize\color{blue}{ ↓ }\) \(\normalsize\color{blue}{ ↓ }\) \(\normalsize\color{blue}{ \rm {Area}=~~~~Width~\times ~Length }\)

OpenStudy (solomonzelman):

See ?

OpenStudy (anonymous):

Oh, so can the width be either x-y, or x+y?

OpenStudy (solomonzelman):

yes, it doesn't matter which one is width and which one is length.

OpenStudy (anonymous):

Thanks soooo much!

OpenStudy (solomonzelman):

Anytime...

OpenStudy (anonymous):

Can I ask you another question please

OpenStudy (solomonzelman):

Yes sure.

OpenStudy (anonymous):

If a rectangle has an area of x^2-2x square units. The lenght would be either x or x-2, right?

OpenStudy (solomonzelman):

`x²-2x` = ` x` × `(x+2)` ↓ ↓ ↓ `AREA` = `Length` × `Width`

OpenStudy (solomonzelman):

Like this.

OpenStudy (solomonzelman):

(For next question, please make a new post)

OpenStudy (anonymous):

Thank you. Does it matter which one is the lenght and which one is the widh?

OpenStudy (anonymous):

Ok

OpenStudy (solomonzelman):

No, it doesn't matter which one is height and which one is length:)